OpenStudy (anonymous):

Please Please Please Help With Question....!!!! Martin Peters earns $6.10 per hour for the first 40 hours each week. He is paid 1 1/2 times his regular pay for all hours over 40. Last week he worked 52 hours.

OpenStudy (anonymous):

Find his gross pay.

OpenStudy (anonymous):

put it back please!! I think that was correct!

OpenStudy (anonymous):

well that was wrong time and a half he get 9.15 for the over time so \[40\times 6.10+12\times 9.15\][

OpenStudy (mayah3):

6.10 X 40 + 12 X 9.15 Hope This Helped .o.

OpenStudy (anonymous):

i get \(\$353.80\) dude needs a new job

OpenStudy (mayah3):

You Just Need To Do The Equations And You'll Get The Answer :']

OpenStudy (anonymous):

Lol! Thank you, That is the correct answer. :) CAn you please help ,e with tow more? PLease PLease!

OpenStudy (anonymous):

I mean two more. :)

OpenStudy (mayah3):

I only knew that on answer, and besides..I have to go soon. I'm sorry I can't help with the rest :(

OpenStudy (mayah3):

Have a good night ^_^

OpenStudy (anonymous):

go ahead and post

OpenStudy (anonymous):

Sorry for the delay just a minute. I will post a new one. :)

OpenStudy (anonymous):

Find his FICA tax at 7.65%.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i guess it is \(.0765\times 358.30\) but i know nothing about fica

OpenStudy (anonymous):

Thank you!! @Mayah3 You to!

OpenStudy (anonymous):

FInd his net pay if other taxes and deductions totaled $73.90.

OpenStudy (anonymous):

you started with $353.80 deducted \(.0765\times 353.80=27.07\) rounded and now deduct another 73.90 in other words compute \[353.80-27.047-73.90\]

OpenStudy (anonymous):

you get 252.85?